# SpeechMapper: Speech-to-text Embedding Projector for LLMs

### v1

@@ 1 Introduction
LLMs have demonstrated unprecedented capabilities on text-based tasks [2, 34, 16], motivating research into extending them to other modalities [31, 9, 13, 14]. For speech, state-of-the-art speech foundation models (SFM) [24, 6] excel at automatic speech recognition (ASR) and speech translation (ST) by leveraging large text decoders that function as internal LMs, but broader spoken language understanding tasks, such as spoken question answering (SQA), require adaptation. To address this gap, recent works have explored integrating SFMs with LLMs. The simplest approach cascades ASR with an LLM, which is effective but discards acoustic cues and adds latency. Similarly, mapping speech into discrete tokens reduces the modality gap but is slow and sensitive to quantization quality [28, 21, 3], underscoring the need for more efficient, content-preserving speech-to-LLM integration. An alternative line of research focuses on directly bridging the SFM encoder outputs to LLMs: early methods inject speech via cross-modal adapters or prepend speech embeddings to the token sequence [25, 11], whereas complex systems jointly instruction-tune (IT) both the SFM encoder and the LLM across multiple tasks [13, 30, 29]. Despite their effectiveness, these methods are resource-intensive, require high-end GPUs and large speech datasets, and the added adapters can lead to prompt overfitting, reducing generalization to unseen tasks.
@@
In contrast to joint SFM–LLM training, recent work maps speech into embeddings that the LLM can consume directly, preserving its original text capabilities. Wav2Prompt [10] trains a speech encoder to generate LLM-like embeddings, enabling zero-shot generalization but still requiring full LLM forward passes during training. SSR [29] similarly learns a speech-to-LLM projector, but through costly distillation with an explicit aligner, followed by LLM fine-tuning. The proposed SpeechMapper follows this line of work. We remove key limitations by avoiding sequence aligners and reducing reliance on costly LLM forward passes, while still producing a projector that generalizes to unseen speech-to-text tasks.
@@
Our SpeechMapper projector maps SFM embeddings directly into the target model’s embedding space through a two-stage training procedure: pretraining and adaptation. Pretraining uses only the LLM’s embedding layer, decoupling training from its costly forward computations, thereby scaling independently of LLM size and enabling the use of more affordable V100 GPUs. Adaptation then rapidly fine-tunes a pretrained SpeechMapper in either a task-agnostic or task-specific manner, keeping the core LLM unchanged. This produces, respectively, a zero-shot speech LLM or a task-specific specialist model, all in just 1.5 h of training on an A100 GPU for an 8–9B LLM. Across ST and SQA experiments, SpeechMapper matches or surpasses the best instruction-following speech LLM from the IWSLT25 short track [1], while requiring far less data and compute, offering a highly resource-efficient path to speech-enabled LLMs.
@@ 2.2 Stage 1 – Pretraining
SpeechMapper is trained using ASR data only. Training minimizes the mean squared error (MSE) loss between the projector’s output embeddings and the corresponding LLM embeddings of the target sentences. A direct MSE application is challenging due to the length mismatch: speech sequences are significantly longer than their text counterparts. Similar works [10, 29] handle this via explicit alignment. Instead, we pad the target embedding sequence with a special pad token to match the projector’s output length. Our assumption is that by forcing some speech vectors to correspond to pad embeddings, the model is forced to shift semantic information toward the beginning of the sequence, implicitly modeling sequence length. Thus, during pretraining, SpeechMapper learns to produce the target embeddings followed by an arbitrary number of pad embeddings. [^2] Successive compression in SpeechMapper keeps the ratio of target to pad embeddings close to or below 1:1. The pad embedding is a reserved LLM token, allowing the LLM to ignore it during inference without additional processing.
In preliminary experiments, we observed that once the first pad token is produced, the model tends to generate subsequent pads easily. To prevent undue emphasis on padding, we split the MSE loss into two parts: a higher-weighted $\text{MSE}_{\text{word}}$ for words and the first pad token, and a lower-weighted $\text{MSE}_{\text{pad}}$ for the remaining pads (Eq. 1), with $\alpha\in{[1,9]}$ controlling the word-to-pad ratio. Masking ensures each loss only considers the relevant embeddings. Following [29], we also incorporate the standard cosine embedding loss to emphasize critical embedding dimensions, which we found to be beneficial. Eq. 2 presents the stage 1 loss, where $\gamma$ controls the cosine weight.
@@ 3 Experimental Settings
Backbones: We use seamless-m4t-v2-large [6] as the SFM, [^4] We also experimented with speech-only pretrained models [5, 7], but stage 1 results were considerably worse. Results are presented in Appendix Section A.2, Table 3. extracting features from the 24th speech encoder layer, and averaging every two frames before feeding them to our projector. We train SpeechMapper for two LLM backbones: Llama-3.1-8B-Instruct [34] and EuroLLM-9B-Instruct [19]. Although stage 1 is free of LLM forward computation and can scale to larger models, we restrict our experiments to 8–9B LLMs because of the many required inferences.
Data: SpeechMapper is pretrained and adapted in task-agnostic settings using the 960 h LibriSpeech (LS) corpus [22]. For task-specific ST/SQA IT, we use the training set used in our baseline [17], sampling a random subset since models train for only 1K steps (considerably less than one epoch). For evaluation, ST experiments use EuroParlST [15] and CoVoST2 [36], while SQA experiments use SpokenSQUAD [18] and LibriSQA [38]. Our prompts are presented in Appendix Section B.3.
@@
Inference and evaluation: LLM inference for the toplines (transcripts/ASR+LLM), BEST-IWSLT25-IF, and our model are performed using the transformers library [37] in greedy mode with a maximum of 150 tokens. We evaluate models on ST and SQA. For ST we compute COMET [^6] Unbabel/wmt22-comet-da [27] using both the reference transcription and translation, and by multiplying scores by 100 for increased readability. We believe COMET is a better metric than BLEU4 for evaluating LLM translations, since BLEU4 heavily penalizes rephrasing. For SQA, we follow the same LLM-as-judge protocol and scripts from BEST-IWSLT25-IF: evaluation is framed as a binary “yes/no” equivalence check using BERGEN [26]. This setup presents the reference context, question, and both generated and reference answers. We report average accuracy across four LLM judges: EuroLLM-9B-Instruct, Gemma3-12B/27B-Instruct [32], and Llama3.1-70B-Instruct. Presented results for stage 2 models correspond to averages obtained by running the IT stage twice using different seeds.
@@ 4.1 Speech Translation
The best zero-shot SpeechMapper (EuroLLM + Stage 2 [ASR CE+MSE]) outperforms Seamless ST across three EuroParl language pairs, despite never being trained on ST data. Our in-domain ST further improves performance by an average of 6.9 COMET points, closing the gap with BEST-IWSLT25-IF and demonstrating the versatility of our task-agnostic pretrained block: by quick in-domain IT, we are able to build a model that rivals strong baselines trained for much longer time and far more data. Lastly, compared to Wav2Prompt’s reported BLEU4 scores on EuroParl en-es and en-fr (13.8/9.2) [10], our approach already substantially outperforms it after stage 1 alone (27.8/24.8). [^9] SacreBLEU (v2.3.1): nrefs:1||case:mixed||eff:no||tok:13a||smooth:exp
@@ A.2 SpeechMapper Stage 1 Hyper-parameters
In this paper we present results using SeamlessM4T-v2-large as SFM. Table 3 presents stage-1 SpeechMappers trained using averaged features extracted from the 10th layer of two SSL-based speech representation models: wav2vec 2.0 [5] and mHuBERT-147 [7]. We observe that SSL models lag behind the model trained using Seamless, especially in out-of-domain settings. We believe this is due to seamless explicitly modeling modality alignment during training [6].
@@ B.1 Other Results
Table 8 present stage 1 and 2 SpeechMapper ASR results over four datasets: LS, VoxPopuli (VP, Wang et al. [35]), CommonVoice v17 (CV, Ardila et al. [4]), and Fleurs [8]. We observe that EuroLLM stage 1 scores are considerably higher than Llama scores, despite both converging in competitive stage 2 models. After qualitative assessment, we believe EuroLLM’s main limitation lies in its verbosity. Despite clear prompt instructions favoring short and direct responses, the model frequently generated multi-line outputs, leading to significant penalties under ASR metrics. In Appendix Section C.1 we present generation examples, highlighting in Table 13 this LLM’s tendency to ignore the prompt. In Appendix Section B.2 we discuss limitations of the WER metric in this particular setting.
@@ B.2 On the limitations of WER/CER for LLM-based ASR evaluation
In this work, due to lack of a better metric, and as in similar bibliography in Speech LLMs [10, 13, 30], we adopt WER/CER as metrics for assessing the capability of our final pipeline to produce transcriptions from corresponding speech. However, we find that this metric overly penalizes our decodings, compared to SFM models such as Whisper or Seamless. In this section we elaborate the reasons, supported by the examples presented in Table 10.

### v2

@@ 1 Introduction
LLMs have demonstrated unprecedented capabilities on text-based tasks [gpt4, llama, jiang2024mixtral], motivating research into extending them to other modalities [team2023gemini, defossez2024moshi, hu2024wavllm, huang2024audiogpt]. For speech, state-of-the-art speech foundation models (SFM) [whisper, barrault2023seamlessm4t] excel at automatic speech recognition (ASR) and speech translation (ST) by leveraging large text decoders that function as internal LMs, but broader spoken language understanding tasks, such as spoken question answering (SQA), require adaptation. To address this gap, recent works have explored integrating SFMs with LLMs. The simplest approach cascades ASR with an LLM, which is effective but discards acoustic cues and adds latency. Similarly, mapping speech into discrete tokens reduces the modality gap but is slow and sensitive to quantization quality [rubenstein2023audiopalm, nguyen2024spirit, ambilduke2025], underscoring the need for more efficient, content-preserving speech-to-LLM integration. An alternative line of research focuses on directly bridging the SFM encoder outputs to LLMs: early methods inject speech via cross-modal adapters or prepend speech embeddings to the token sequence [whispering-llama, Fathullah2023PromptingLL], whereas complex systems jointly instruction-tune (IT) both the SFM encoder and the LLM across multiple tasks [hu2024wavllm, salmonn, tan2024ssr]. Despite their effectiveness, these methods are resource-intensive, require high-end GPUs and large speech datasets, and the added adapters can lead to prompt overfitting, reducing generalization to unseen tasks.
@@
In contrast to joint SFM–LLM training, recent work maps speech into embeddings that the LLM can consume directly, preserving its original text capabilities. Wav2Prompt [wav2Prompt] trains a speech encoder to generate LLM-like embeddings, enabling zero-shot generalization but still requiring full LLM forward passes during training. SSR [tan2024ssr] similarly learns a speech-to-LLM projector, but through costly distillation with an explicit aligner, followed by LLM fine-tuning. The proposed SpeechMapper follows this line of work. We remove key limitations by avoiding sequence aligners and reducing reliance on costly LLM forward passes, while still producing a projector that generalizes to unseen speech-to-text tasks.
@@
Our SpeechMapper projector maps SFM embeddings directly into the target model’s embedding space through a two-stage training procedure: pretraining and adaptation. Pretraining uses only the LLM’s embedding layer, decoupling training from its costly forward computations, thereby scaling independently of LLM size and enabling the use of more affordable V100 GPUs. Adaptation then rapidly fine-tunes a pretrained SpeechMapper in either a task-agnostic or task-specific manner, keeping the core LLM unchanged. This produces, respectively, a zero-shot speech LLM or a task-specific specialist model, all in just 1.5 h of training on an A100 GPU for an 8–9B LLM. Across ST and SQA experiments, SpeechMapper matches or surpasses the best instruction-following speech LLM from the IWSLT25 short track [agostinelli-etal-2025-findings], while requiring far less data and compute, offering a highly resource-efficient path to speech-enabled LLMs.
@@ 2.2 Stage 1 – Pretraining
SpeechMapper is trained using ASR data only. Training minimizes the mean squared error (MSE) loss between the projector’s output embeddings and the corresponding LLM embeddings of the target sentences. A direct MSE application is challenging due to the length mismatch: speech sequences are significantly longer than their text counterparts. Similar works [wav2Prompt, tan2024ssr] handle this via explicit alignment. Instead, we pad the target embedding sequence with a special pad token to match the projector’s output length. Our assumption is that by forcing some speech vectors to correspond to pad embeddings, the model is forced to shift semantic information toward the beginning of the sequence, implicitly modeling sequence length. Thus, during pretraining, SpeechMapper learns to produce the target embeddings followed by an arbitrary number of pad embeddings. [^2] Successive compression in SpeechMapper keeps the ratio of target to pad embeddings close to or below 1:1. The pad embedding is a reserved LLM token, allowing the LLM to ignore it during inference without additional processing.
In preliminary experiments, we observed that once the first pad token is produced, the model tends to generate subsequent pads easily. To prevent undue emphasis on padding, we split the MSE loss into two parts: a higher-weighted $\text{MSE}_{\text{word}}$ for words and the first pad token, and a lower-weighted $\text{MSE}_{\text{pad}}$ for the remaining pads (Eq. 1), with $\alpha\in{[1,9]}$ controlling the word-to-pad ratio. Masking ensures each loss only considers the relevant embeddings. Following [tan2024ssr], we also incorporate the standard cosine embedding loss to emphasize critical embedding dimensions, which we found to be beneficial. Eq. 2 presents the stage 1 loss, where $\gamma$ controls the cosine weight.
@@ 3 Experimental Settings
Backbones: We use seamless-m4t-v2-large [barrault2023seamlessm4t] as the SFM, [^4] We also experimented with speech-only pretrained models [NEURIPS2020_92d1e1eb, boito2024mhubert], but stage 1 results were considerably worse. Results are presented in Appendix Section A.2, Table 3. extracting features from the 24th speech encoder layer, and averaging every two frames before feeding them to our projector. We train SpeechMapper for two LLM backbones: Llama-3.1-8B-Instruct [llama] and EuroLLM-9B-Instruct [martins2024eurollm]. Although stage 1 is free of LLM forward computation and can scale to larger models, we restrict our experiments to 8–9B LLMs because of the many required inferences.
Data: SpeechMapper is pretrained and adapted in task-agnostic settings using the 960 h LibriSpeech (LS) corpus [panayotov2015librispeech]. For task-specific ST/SQA IT, we use the training set used in our baseline [lee-etal-2025-naver], sampling a random subset since models train for only 1K steps (considerably less than one epoch). For evaluation, ST experiments use EuroParlST [iranzo2020europarl] and CoVoST2 [wang2020covost], while SQA experiments use SpokenSQUAD [lee2018spoken] and LibriSQA [zhao2024librisqa]. Our prompts are presented in Appendix Section B.3.
@@
Inference and evaluation: LLM inference for the toplines (transcripts/ASR+LLM), BEST-IWSLT25-IF, and our model are performed using the transformers library [wolf2019huggingface] in greedy mode with a maximum of 150 tokens. We evaluate models on ST and SQA. For ST we compute COMET [^6] Unbabel/wmt22-comet-da [rei-etal-2022-comet] using both the reference transcription and translation, and by multiplying scores by 100 for increased readability. We believe COMET is a better metric than BLEU4 for evaluating LLM translations, since BLEU4 heavily penalizes rephrasing. For SQA, we follow the same LLM-as-judge protocol and scripts from BEST-IWSLT25-IF: evaluation is framed as a binary “yes/no” equivalence check using BERGEN [rau-etal-2024-bergen]. This setup presents the reference context, question, and both generated and reference answers. We report average accuracy across four LLM judges: EuroLLM-9B-Instruct, Gemma3-12B/27B-Instruct [gemma3], and Llama3.1-70B-Instruct. Presented results for stage 2 models correspond to averages obtained by running the IT stage twice using different seeds.
@@ 4.1 Speech Translation
The best zero-shot SpeechMapper (EuroLLM + Stage 2 [ASR CE+MSE]) outperforms Seamless ST across three EuroParl language pairs, despite never being trained on ST data. Our in-domain ST further improves performance by an average of 6.9 COMET points, closing the gap with BEST-IWSLT25-IF and demonstrating the versatility of our task-agnostic pretrained block: by quick in-domain IT, we are able to build a model that rivals strong baselines trained for much longer time and far more data. Lastly, compared to Wav2Prompt’s reported BLEU4 scores on EuroParl en-es and en-fr (13.8/9.2) [wav2Prompt], our approach already substantially outperforms it after stage 1 alone (27.8/24.8). [^9] SacreBLEU (v2.3.1): nrefs:1||case:mixed||eff:no||tok:13a||smooth:exp
@@ A.2 SpeechMapper Stage 1 Hyper-parameters
In this paper we present results using SeamlessM4T-v2-large as SFM. Table 3 presents stage-1 SpeechMappers trained using averaged features extracted from the 10th layer of two SSL-based speech representation models: wav2vec 2.0 [NEURIPS2020_92d1e1eb] and mHuBERT-147 [boito2024mhubert]. We observe that SSL models lag behind the model trained using Seamless, especially in out-of-domain settings. We believe this is due to seamless explicitly modeling modality alignment during training [barrault2023seamlessm4t].
@@ B.1 Other Results
Table 8 present stage 1 and 2 SpeechMapper ASR results over four datasets: LS, VoxPopuli (VP, voxpopuli), CommonVoice v17 (CV, commonvoice:17), and Fleurs [conneau2023fleurs]. We observe that EuroLLM stage 1 scores are considerably higher than Llama scores, despite both converging in competitive stage 2 models. After qualitative assessment, we believe EuroLLM’s main limitation lies in its verbosity. Despite clear prompt instructions favoring short and direct responses, the model frequently generated multi-line outputs, leading to significant penalties under ASR metrics. In Appendix Section C.1 we present generation examples, highlighting in Table 13 this LLM’s tendency to ignore the prompt. In Appendix Section B.2 we discuss limitations of the WER metric in this particular setting.
@@ B.2 On the limitations of WER/CER for LLM-based ASR evaluation
In this work, due to lack of a better metric, and as in similar bibliography in Speech LLMs [wav2Prompt, hu2024wavllm, salmonn], we adopt WER/CER as metrics for assessing the capability of our final pipeline to produce transcriptions from corresponding speech. However, we find that this metric overly penalizes our decodings, compared to SFM models such as Whisper or Seamless. In this section we elaborate the reasons, supported by the examples presented in Table 10.